# Chow groups of one-dimensional noetherian domains

### v2

@@ 2. Divisors and Chow groups
Corollary 2.8 shows that $\varphi$ is a well-defined epimorphism. The subgroup $R$ is in its kernel since $f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\mathfrak{Q}_{i}))=g_{i}\mathfrak{p}_{i}$. Conversely, suppose that $(\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i},[\mathfrak{A}])$ is in the kernel of $\varphi$. Then $\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i}$ is in the image of $f_{*}$. By Corollary 2.8 this implies $g_{i}\mid a_{i}$ for all $i$. So $(0,[\mathfrak{A}\prod_{i=1}^{r}\mathfrak{Q}_{i}^{a_{i}/g_{i}}])$ is also in the kernel of $\varphi$. By Remark 2.10 this is only possible if $\mathfrak{A}\prod_{i=1}^{r}\mathfrak{Q}_{i}^{a_{i}/g_{i}}\in N$, i.e. $[\mathfrak{A}\prod_{i=1}^{r}\mathfrak{Q}_{i}^{a_{i}/g_{i}}]\in G/N$ is trivial. Hence $R$ is the kernel of $\varphi$. ∎
@@
    Using Lemma 2.1, find $\mathfrak{A}\in\mathcal{I}({\tilde{\mathcal{O}}})$ such that $f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\mathfrak{A}))=D$.
@@
    Compute the class group $\operatorname{Pic}({\tilde{\mathcal{O}}})$.
@@
    From the factorization of the conductor $\mathcal{F}$ in ${\tilde{\mathcal{O}}}$ obtain the non-invertible maximal ideals of $\mathcal{O}$ and generators of $\ker(f_{*})$, cf. Remark 2.10.
@@
    If $[\mathfrak{A}]\notin\overline{\operatorname{div}}_{\tilde{\mathcal{O}}}^{-1}(\ker(\overline{f_{*}}))$, return false.
@@
    Pick a $\mathfrak{B}\in\operatorname{div}_{\tilde{\mathcal{O}}}^{-1}(\ker(f_{*}))$ such that $\mathfrak{A}\mathfrak{B}=\alpha{\tilde{\mathcal{O}}}$ is principal.
@@
Suppose first $D=\operatorname{div}_{\mathcal{O}}(\beta)$ is a principal divisor. Remark 2.7 shows that $D=f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\beta{\tilde{\mathcal{O}}}))\in\operatorname{im}(f_{*})$. The choice of $\mathfrak{A}$ yields $D=f_{*}(\operatorname{div}(\mathfrak{A}))$ and thus $\beta^{-1}\mathfrak{A}\in\operatorname{div}^{-1}(\ker(f_{*}))$. Hence there exists some ideal $\mathfrak{B}\in\operatorname{div}_{\tilde{\mathcal{O}}}^{-1}(\ker(f_{*}))$ such that $\mathfrak{A}\mathfrak{B}=\alpha{\tilde{\mathcal{O}}}$ is principal. Again by Remark 2.7 it follows that
$\operatorname{div}_{\mathcal{O}}(\alpha)=f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\mathfrak{A}))+f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\mathfrak{B}))=D+0=D\>.$
@@
Provided that we can compute $\operatorname{Pic}({\tilde{\mathcal{O}}})$ and the conductor $\mathcal{F}$ as well as factor $\mathcal{F}$ into prime ideals, Theorem 2.11 yields an effective method to compute $\operatorname{Chow}(\mathcal{O})$. To be able to run Algorithm 2.12, we also need to check if a fractional ideal of ${\tilde{\mathcal{O}}}$ is principal and if so, compute a generator. For orders $\mathcal{O}$ in a global field $K$, there are well known algorithms to solve these tasks.
We end this section by recalling the definition of Picard groups and discuss the canonical homomorphism $\operatorname{Pic}(\mathcal{O})\to\operatorname{Chow}(\mathcal{O})$.
is called the Picard group of $\mathcal{O}$. Being invertible is a local property. For an invertible fractional ideal $\mathfrak{a}$ of $\mathcal{O}$ and $\mathfrak{p}\in\mathbb{P}(\mathcal{O})$ there exists some $a_{\mathfrak{p}}\in K^{*}$ such that $\mathfrak{a}_{\mathfrak{p}}=a\mathcal{O}_{\mathfrak{p}}$, see [8, Satz (12.4)] for details. Hence we can define
which induces a homomorphism
The homomorphism $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}}),\mathfrak{a}\mapsto\mathfrak{a}{\tilde{\mathcal{O}}}$ maps principal ideals to principal ideals and therefore it induces a canonical homomorphism
@@
Since every element of $\operatorname{Pic}({\tilde{\mathcal{O}}})$ is represented by a product of prime ideals of ${\tilde{\mathcal{O}}}$ coprime to $\mathcal{F}$, the canonical homomorphism $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$ is surjective, see also [8, p. 79].
Lemma 2.1 shows that the homomorphism $\overline{\operatorname{div}}_{\mathcal{O}}\colon\operatorname{Pic}(\mathcal{O})\to\operatorname{Chow}(\mathcal{O})$ is the composition of the canonical homomorphisms
@@
Since $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$ is surjective, the map $\overline{\operatorname{div}}_{\mathcal{O}}$ is surjective if and only if $\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$ is. By Corollary 2.8 that holds if and only if $\overline{f_{*}}$ is surjective. Theorem 2.9 shows that this is equivalent to all local Chow groups being trivial.
Similarly, $\overline{\operatorname{div}}_{\mathcal{O}}$ is injective if and only if $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$ and $\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$ are. So the second assertion follows again from Corollary 2.8. ∎
@@ 3. Orders in number fields with trivial Chow groups
In this section we show that number fields often contain orders with trivial Chow group. So let $K$ be an algebraic number field with maximal order ${\tilde{\mathcal{O}}}$. We start with recalling Furtwängler’s classification of conductor ideals.
Definition 3.1.
An ideal $\mathfrak{A}$ of ${\tilde{\mathcal{O}}}$ is called a conductor ideal if there exists some order in $K$ with conductor $\mathfrak{A}$.
Theorem 3.2 (Furtwängler).
@@
    Let $\mathfrak{A}$ and $\mathfrak{B}$ be ideals of ${\tilde{\mathcal{O}}}$ with coprime norms. Then $\mathfrak{A}\mathfrak{B}$ is a conductor ideal if and only if $\mathfrak{A}$ and $\mathfrak{B}$ are conductor ideals.
@@
    Let $p$ be a prime number and let $p{\tilde{\mathcal{O}}}=\mathfrak{P}_{1}^{e_{1}}\cdot\ldots\cdot\mathfrak{P}_{r}^{e_{r}}$ be the factorization into prime ideals of ${\tilde{\mathcal{O}}}$. Then $\mathfrak{P}_{1}^{k_{1}}\cdot\ldots\cdot\mathfrak{P}_{r}^{k_{r}}$ is a conductor ideal if and only if for all $1\leq i\leq r$ such that ${\tilde{\mathcal{O}}}/\mathfrak{P}_{i}\cong\mathbb{F}_{p}$ and $k_{i}\equiv 1\pmod{e_{i}}$ there exists some index $j\neq i$ such that $k_{j}>\frac{k_{i}-1}{e_{i}}e_{j}$.
@@
See for example [2, 3, 6]. ∎
Corollary 3.3.
If $K\neq\mathbb{Q}$ and $a\in\mathbb{Z}$ is non-zero, then $a{\tilde{\mathcal{O}}}$ is a conductor ideal.
In the following results, we use orders that arise from ideals of ${\tilde{\mathcal{O}}}$. This is explained in the following lemma.
Lemma 3.4.
Let $\mathfrak{A}$ be an ideal of ${\tilde{\mathcal{O}}}$ and let $p$ be a prime with $\mathbb{Z}\cap\mathfrak{A}\subseteq p\mathbb{Z}$.
  1. (1)
    $\mathbb{Z}+\mathfrak{A}$ is an order in $K$.
  2. (2)
    $\mathfrak{A}$ is a conductor ideal if and only if $\mathbb{Z}+\mathfrak{A}$ has conductor $\mathfrak{A}$ in which case $\mathbb{Z}+\mathfrak{A}$ is the smallest order of conductor $\mathfrak{A}$ in $K$.
  3. (3)
    $\mathfrak{p}:=p\mathbb{Z}+\mathfrak{A}$ is a maximal ideal of $\mathbb{Z}+\mathfrak{A}$ over $p$ and $(\mathbb{Z}+\mathfrak{A})/\mathfrak{p}\cong\mathbb{F}_{p}$.
  4. (4)
    If $\mathfrak{A}$ is a conductor ideal, then $\mathfrak{p}$ is the unique non-invertible maximal ideal of $\mathbb{Z}+\mathfrak{A}$ over $p$.
Proof.
The $\mathbb{Z}$-module $\mathbb{Z}+\mathfrak{A}$ is finitely generated and a subring of $K$, hence an order. Suppose $\mathcal{O}^{\prime}$ is an order in $K$ of conductor $\mathfrak{A}$. The inclusions
$\mathfrak{A}\subseteq\mathbb{Z}+\mathfrak{A}\subseteq\mathcal{O}^{\prime}$
show that $\mathbb{Z}+\mathfrak{A}$ also has conductor $\mathfrak{A}$.
From $\mathfrak{p}=p(\mathbb{Z}+\mathfrak{A})+\mathfrak{A}$ it is clear that $\mathfrak{p}$ is an ideal of $\mathbb{Z}+\mathfrak{A}$ and $[\mathbb{Z}+\mathfrak{A}:\mathfrak{p}]=p$ shows that it is indeed maximal. Suppose now $\mathfrak{A}$ is a conductor ideal. Then $\mathbb{Z}+\mathfrak{A}$ has conductor $\mathfrak{A}$ and $\mathfrak{p}+\mathfrak{A}=\mathfrak{p}$ implies that $\mathfrak{p}$ is not invertible. Let $\mathfrak{q}$ be another non-invertible maximal ideal of $\mathbb{Z}+\mathfrak{A}$ over $p$. Then $\mathfrak{q}+\mathfrak{A}=\mathfrak{q}$ and
$\mathfrak{p}^{2}=p^{2}\mathbb{Z}+p\mathfrak{A}+\mathfrak{A}^{2}\subseteq\mathfrak{q}+\mathfrak{A}=\mathfrak{q}$
@@
We remark that two different orders in $K$ might have the same conductor, see Example 4.2.
@@
By [5, Theorem 1] there are prime numbers $p_{1},\dots,p_{r}$ such that the maximal ideals of ${\tilde{\mathcal{O}}}$ of (absolute) degree $1$ over these primes generate the class group $\operatorname{Pic}({\tilde{\mathcal{O}}})$. By Corollary 3.3, the order $\mathcal{O}=\mathbb{Z}+(p_{1}\cdot\ldots\cdot p_{r}){\tilde{\mathcal{O}}}$ has conductor $(p_{1}\cdot\ldots\cdot p_{r}){\tilde{\mathcal{O}}}$. For $1\leq i\leq r$ there exists by Lemma 3.4 a unique maximal ideal $\mathfrak{p}_{i}$ of $\mathcal{O}$ over $p_{i}$. Theorem 2.9 therefore implies that $\operatorname{Chow}({\tilde{\mathcal{O}}}_{\mathfrak{p}_{i}})$ is trivial and thus $\overline{f_{*}}$ is onto.
Let $\mathfrak{P}$ be a maximal ideal of ${\tilde{\mathcal{O}}}$ over $p_{i}$ of degree $1$. Then $\mathfrak{P}\cap L$ is principal, generated by $a\in L^{*}$ say. So the integral element $a\in{\tilde{\mathcal{O}}}$ has absolute norm $p^{m}$. Hence $\operatorname{ord}_{\mathfrak{p}_{i}}(a)=m$ and $\operatorname{ord}_{\mathfrak{q}}(a)=0$ for all $\mathfrak{q}\in\mathbb{P}(\mathcal{O})$ different from $\mathfrak{p}_{i}$. Thus
@@
and $[\mathfrak{P}]^{m}=[\mathfrak{P}^{m}/a]$ lies in the kernel of $\operatorname{Pic}({\tilde{\mathcal{O}}})\cong\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$. Since $m$ is coprime to the class number of ${\tilde{\mathcal{O}}}$, we conclude that $[\mathfrak{P}]$ itself lies in the kernel. Since the maximal ideals of ${\tilde{\mathcal{O}}}$ of degree 1 over the primes $p_{1},\dots,p_{r}$ generate the class group of ${\tilde{\mathcal{O}}}$, we conclude that $\operatorname{Chow}(\mathcal{O})=\operatorname{im}(\overline{f_{*}})$ is trivial. ∎
@@
In case we have found an order $\mathcal{O}$ with trivial Chow group it is easy to find infinitely many suborders which have trivial Chow group as well. Indeed, intersecting $\mathcal{O}$ with the order $\mathbb{Z}+p{\tilde{\mathcal{O}}}$ for a completely splitting prime $p$ coprime to the index $[{\tilde{\mathcal{O}}}:\mathcal{O}]$ yields an order which has trivial Chow group as well.

### v3

@@ 2. Divisors and Chow groups
(1) The given generators lie in $\ker(f_{*})$. Conversely, let $D:=\sum_{i,j}a_{i,j}\mathfrak{P}_{i,j}\in\ker(f_{*})$. Then $\sum_{j}a_{i,j}d_{i,j}=0$ for all $i$, see Lemma 2.5. Therefore
$D+\sum_{i,j}a_{i,j}(\tfrac{d_{i,j}}{g_{i}}Q_{i}-\mathfrak{P}_{i,j})=\sum_{i,j}\tfrac{a_{i,j}d_{i,j}}{g_{i}}Q_{i}=\sum_{i}\tfrac{1}{g_{i}}(\sum_{j}a_{i,j}d_{i,j})Q_{i}=0\>.$
The result follows.
(2) Lemma 2.5 shows that $\varphi$ is a well-defined epimorphism. The subgroup $R$ is in its kernel since $f_{*}(Q_{i})=g_{i}\mathfrak{p}_{i}$. Conversely, suppose that $(\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i},[A])$ is in the kernel of $\varphi$. Then $\sum_{i=1}^{r}a_{i}\mathfrak{p}_{i}$ is in the image of $f_{*}$. By Lemma 2.5 this implies $g_{i}\mid a_{i}$ for all $i$. So $(0,[A+\sum_{i=1}^{r}\frac{a_{i}}{g_{i}}Q_{i}])$ is also in the kernel of $\varphi$. By (1) this is only possible if $A+\sum_{i=1}^{r}\frac{a_{i}}{g_{i}}Q_{i}\in\ker(f_{*})$, i.e. $[A+\sum_{i=1}^{r}\frac{a_{i}}{g_{i}}Q_{i}]\in G/\ker(\overline{f_{*}})$ is trivial. Hence $R$ is the kernel of $\varphi$. ∎
@@
    Using Proposition 2.2, find $A\in\operatorname{Div}({\tilde{\mathcal{O}}})$ such that $f_{*}(A)=D$.
@@
    Compute $\operatorname{Chow}({\tilde{\mathcal{O}}})\cong\operatorname{Pic}({\tilde{\mathcal{O}}})$.
@@
    From the factorization of the conductor $\mathcal{F}$ in ${\tilde{\mathcal{O}}}$ obtain the non-invertible maximal ideals of $\mathcal{O}$ and generators of $\ker(f_{*})$, cf. Theorem 2.7.
@@
    If $[A]\notin\ker(\overline{f_{*}})$, return false.
@@
    Pick some $B\in\ker(f_{*})$ such that $A+B=\operatorname{div}_{\mathcal{O}}(\alpha)$ is principal.
@@
Suppose first $D=\operatorname{div}_{\mathcal{O}}(\beta)$ is a principal divisor. Equation (2.1) shows that $D=f_{*}(\operatorname{div}_{\tilde{\mathcal{O}}}(\beta))\in\operatorname{im}(f_{*})$. The choice of $A$ yields $D=f_{*}(A)$ and thus $A-\operatorname{div}_{\tilde{\mathcal{O}}}(\beta)\in\ker(f_{*})$. Hence there exists some $B\in\ker(f_{*})$ such that $A+B$ is principal. So the element $\alpha$ from step (6) of the algorithm satisfies
$\operatorname{div}_{\mathcal{O}}(\alpha)=f_{*}(A)+f_{*}(B)=D+0=D\>.$
@@
Provided that we can compute $\operatorname{Chow}({\tilde{\mathcal{O}}})\cong\operatorname{Pic}({\tilde{\mathcal{O}}})$ and the conductor $\mathcal{F}$ as well as factor $\mathcal{F}$ into prime ideals, Theorem 2.7 yields an effective method to compute $\operatorname{Chow}(\mathcal{O})$. To be able to run Algorithm 2.8, we also need to check if a given divisor of ${\tilde{\mathcal{O}}}$ (or equivalently a given fractional ideal of ${\tilde{\mathcal{O}}}$) is principal and if so, compute a generator. For orders $\mathcal{O}$ in a global field $K$, there are well known algorithms to solve these tasks.
We close this section by discussing the canonical homomorphism $\operatorname{Pic}(\mathcal{O})\to\operatorname{Chow}(\mathcal{O})$. The homomorphism $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}}),\mathfrak{a}\mapsto\mathfrak{a}{\tilde{\mathcal{O}}}$ maps principal ideals to principal ideals and therefore it induces a canonical homomorphism
@@
Since every element of $\operatorname{Pic}({\tilde{\mathcal{O}}})$ is represented by a product of prime ideals of ${\tilde{\mathcal{O}}}$ coprime to $\mathcal{F}$, the canonical homomorphism $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$ is surjective, see also [9, p. 79].
Proposition 2.2 shows that the homomorphism $\overline{\operatorname{div}}_{\mathcal{O}}\colon\operatorname{Pic}(\mathcal{O})\to\operatorname{Chow}(\mathcal{O})$ is the composition of the canonical homomorphisms
@@
Since $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$ is surjective, the map $\overline{\operatorname{div}}_{\mathcal{O}}$ is surjective if and only if $\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$ is. By Lemma 2.5 that holds if and only if $\overline{f_{*}}$ is surjective. Theorem 2.6 shows that this is equivalent to all local Chow groups being trivial.
Similarly, $\overline{\operatorname{div}}_{\mathcal{O}}$ is injective if and only if $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$ and $\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$ are. So the second assertion follows again from Lemma 2.5. ∎
@@ 3. Orders in number fields with trivial Chow groups
We are going to show that number fields often contain orders with trivial Chow group. We start with a more general result which might be of independent interest, but we will only use it in the case that $R=\mathbb{Z}$ and ${\tilde{\mathcal{O}}}$ is the maximal order of an algebraic number field. In this case, part (2) can also be deduced from Furtwängler’s classification of conductors, see [2, 3, 7].
Lemma 3.1.
Let $R\subsetneq{\tilde{\mathcal{O}}}$ be a finite extension of Dedekind rings. Let $\mathfrak{a}$ be a non-zero ideal of $R$ and set $\mathcal{O}=R+\mathfrak{a}{\tilde{\mathcal{O}}}$.
@@
    $\mathcal{O}$ is an $R$-subalgebra of ${\tilde{\mathcal{O}}}$ which is finitely generated over $R$ and has the same field of fractions as ${\tilde{\mathcal{O}}}$.
@@
    The conductor of $\mathcal{O}$ is $\mathfrak{a}{\tilde{\mathcal{O}}}$.
  3. (3)
    Let $\mathfrak{p}\in\mathbb{P}(R)$ with $\mathfrak{a}\subseteq\mathfrak{p}$. Then $\mathfrak{P}:=\mathfrak{p}+\mathfrak{a}{\tilde{\mathcal{O}}}$ is the unique maximal ideal of $\mathcal{O}$ over $\mathfrak{p}$. It is non-invertible and $\mathcal{O}/\mathfrak{P}\cong R/\mathfrak{p}$.
@@
Part (1) is clear.
(2) Let $\mathcal{F}$ be the conductor of $\mathcal{O}$. Then $\mathfrak{a}{\tilde{\mathcal{O}}}\subseteq\mathcal{F}$. To show that $\mathfrak{a}{\tilde{\mathcal{O}}}=\mathcal{F}$, we may pass to the localizations of $R$ and assume that $R$ is local. Then $\mathfrak{a}=fR$ and $\mathcal{F}\cap R=gR$ for some $f,g\in R$. From $f\in\mathcal{F}\cap R=gR$ we find $r\in R$ such that $f=rg$. Let $x\in\mathcal{F}$ and write $x=s+fy$ with $s\in R$ and $y\in{\tilde{\mathcal{O}}}$. Then $s\in R\cap\mathcal{F}=gR$ and $x\in gR+f{\tilde{\mathcal{O}}}\subseteq g{\tilde{\mathcal{O}}}$. This shows $\mathcal{F}=g{\tilde{\mathcal{O}}}$. The inclusions
$\mathcal{O}=R+rg{\tilde{\mathcal{O}}}=R+\mathfrak{a}{\tilde{\mathcal{O}}}\subseteq R+\mathcal{F}=R+g{\tilde{\mathcal{O}}}\subseteq\mathcal{O}$
show $R+rg{\tilde{\mathcal{O}}}=R+g{\tilde{\mathcal{O}}}$ and thus
$r\cdot\mathcal{O}/R=r\cdot(R+g{\tilde{\mathcal{O}}})/R=(R+rg{\tilde{\mathcal{O}}})/R=\mathcal{O}/R\>.$
If $r$ was not a unit, Nakayama’s lemma gives $\mathcal{O}/R=0$ and thus $R=\mathcal{O}$. This case is excluded. So $r\in R^{*}$ and thus $\mathcal{F}=g{\tilde{\mathcal{O}}}=f{\tilde{\mathcal{O}}}=\mathfrak{a}{\tilde{\mathcal{O}}}$ as claimed.
(3) $\mathfrak{P}$ is a proper ideal of $\mathcal{O}$. Thus the inclusions $\mathfrak{p}\subseteq R\cap\mathfrak{P}\subsetneq R$ imply $\mathfrak{p}=\mathfrak{P}\cap R$. So the epimorphism $R\to\mathcal{O}/\mathfrak{P}$ has kernel $\mathfrak{p}$ and therefore $\mathcal{O}/\mathfrak{P}\cong R/\mathfrak{p}$. Thus $\mathfrak{P}$ is maximal and it is non-invertible as $\mathfrak{P}+\mathcal{F}=\mathfrak{p}+\mathfrak{a}{\tilde{\mathcal{O}}}=\mathfrak{P}$. Let $\mathfrak{Q}$ be a maximal ideal of $\mathcal{O}$ over $\mathfrak{p}$. Then
$\mathfrak{P}^{2}=\mathfrak{p}^{2}+\mathfrak{p}\mathfrak{a}{\tilde{\mathcal{O}}}+\mathfrak{a}^{2}{\tilde{\mathcal{O}}}=\mathfrak{p}^{2}+\mathfrak{p}\mathcal{F}\subseteq\mathfrak{p}\mathcal{O}\subseteq\mathfrak{Q}$
@@
For the remainder of this section, let $K$ be an algebraic number field with ring of integers ${\tilde{\mathcal{O}}}$. We will use the previous result with $R=\mathbb{Z}$ to construct orders in $K$ with trivial Chow groups.
@@
We may assume that $K\neq\mathbb{Q}$. By [6, Theorem 1] the class group $\operatorname{Pic}({\tilde{\mathcal{O}}})$ can be generated by a finite set $\mathcal{P}$ of maximal ideals of ${\tilde{\mathcal{O}}}$ of (absolute) degree $1$. Let $p_{1},\dots,p_{r}$ be the different prime numbers below the ideals in $\mathcal{P}$. By Lemma 3.1, the order $\mathcal{O}=\mathbb{Z}+(p_{1}\cdot\ldots\cdot p_{r}){\tilde{\mathcal{O}}}$ has conductor $(p_{1}\cdot\ldots\cdot p_{r}){\tilde{\mathcal{O}}}$ and there exists a unique maximal ideal $\mathfrak{p}_{i}$ of $\mathcal{O}$ over $p_{i}$. Theorem 2.6 therefore implies that $\operatorname{Chow}({\tilde{\mathcal{O}}}_{\mathfrak{p}_{i}})$ is trivial and thus $\overline{f_{*}}\colon\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$ is onto.
Let ${\tilde{\mathcal{O}}}_{L}$ be the ring of integers of $L$ and let $\mathfrak{P}$ be a maximal ideal of ${\tilde{\mathcal{O}}}$ over $p_{i}$ of degree $1$. Then $\mathfrak{P}\cap{\tilde{\mathcal{O}}}_{L}$ is principal, generated by $a\in L^{*}$ say. So we have the following inclusions
${\mathbb{Z}}$${{\tilde{\mathcal{O}}}_{L}}$${{\tilde{\mathcal{O}}}}$${p_{i}\mathbb{Z}}$${a{\tilde{\mathcal{O}}}_{L}}$${\mathfrak{P}}$
Since $a{\tilde{\mathcal{O}}}_{L}=\mathfrak{P}\cap{\tilde{\mathcal{O}}}_{L}$ has degree $1$, the integral element $a\in{\tilde{\mathcal{O}}}$ has absolute norm $p_{i}^{m}$. Hence $a\in K$ satisfies $\operatorname{ord}_{\mathfrak{p}_{i}}(a)=m$ and $\operatorname{ord}_{\mathfrak{q}}(a)=0$ for all $\mathfrak{q}\in\mathbb{P}(\mathcal{O})$ different from $\mathfrak{p}_{i}$. Thus
@@
and $[\mathfrak{P}]^{m}=[\mathfrak{P}^{m}/a]$ lies in the kernel of $\operatorname{Pic}({\tilde{\mathcal{O}}})\cong\operatorname{Chow}({\tilde{\mathcal{O}}})\to\operatorname{Chow}(\mathcal{O})$. Since $m$ is coprime to the class number of ${\tilde{\mathcal{O}}}$, we conclude that $[\mathfrak{P}]$ itself lies in the kernel. Since the maximal ideals of ${\tilde{\mathcal{O}}}$ of degree 1 over the primes $p_{1},\dots,p_{r}$ generate the class group of ${\tilde{\mathcal{O}}}$, Theorem 2.6 shows that $\operatorname{Chow}(\mathcal{O})=\operatorname{im}(\overline{f_{*}})$ is trivial. ∎
@@
We close this section by showing that the number of orders in $K$ with trivial Chow group is either zero or infinite.
Lemma 3.8.
If $K$ contains an order with trivial Chow group, then there exist infinitely many such orders in $K$.
Proof.
Let $\mathcal{O}^{\prime}$ be an order in $K$ has trivial Chow group and set $m=[{\tilde{\mathcal{O}}}:\mathcal{O}^{\prime}]$. By Chebotarev’s Density Theorem, there exist infinitely many primes $p$ coprime to $m$ which completely split in $K$. Hence it suffices to show that $\mathcal{O}:=\mathcal{O}^{\prime}\cap(\mathbb{Z}+p{\tilde{\mathcal{O}}})$ has trivial Chow group. Let $q$ be a prime divisor of $m$ and let $\mathfrak{q}$ be a prime ideal of $\mathcal{O}$ over $q$. Localizing at $\mathfrak{q}$ shows that there exists a unique prime ideal $\mathfrak{Q}$ of $\mathcal{O}^{\prime}$ over $\mathfrak{q}$ and $\mathcal{O}/\mathfrak{q}\cong\mathcal{O}^{\prime}/\mathfrak{Q}$. By Lemma 3.1 the order $\mathbb{Z}+p{\tilde{\mathcal{O}}}$ only has one prime ideal over $p$ and it is of degree $1$. The same must hold for the suborder $\mathcal{O}$. Lemma 2.5 shows that the induced morphism $\overline{f_{*}}\colon\operatorname{Chow}(\mathcal{O}^{\prime})\to\operatorname{Chow}(\mathcal{O})$ is surjective. Hence $\operatorname{Chow}(\mathcal{O})$ is trivial. ∎